# GraphRAFT: Retrieval Augmented Fine-Tuning for Knowledge Graphs on Graph Databases

### v1

@@ 4.3 Finetuning LLM as local subgraph reasoner
Given any question for KG $G$, our trained LLM $L_{1}$ produces guaranteed executable and grounded Cypher queries. If we perform beam search with width $m$, we are able to obtain $m$ valid queries with highest total probabilties. For the simplest questions and sparse $G$, the subgraph produced by the best query may be exactly the answer nodes. However, for more difficult multi-hop questions on dense $G$, the subgraph still contains other relevant nodes (and edges) that are not the answers. In order to obtain only the answer nodes by reasoning over the small subgraph, which often requires reasoning on the textual properties beyond graph patterns, we train an LLM $L_{2}$ to perform the task. We construct the subgraph by executing $m$ queries until it contains some threshold on the size of graph or tokenization required to encode the graph. The outputs of $L_{2}$ can be viewed as selecting or reranking the nodes in the input textualised graph. An example prompt we used provided in Figure 4.
@@ 5 Experiments
We use Neo4j as the graph database. The default database configuration is used. For the main result, we using OpenAI text-embedding-ada-002[61] as the text embedder $LM$ in Equation 1, OpenAI gpt-4o-mini as LLM $L_{0}$ for few-shot entity resolution., gemma2-9b-text2cypher[62, 63] as our LLM $L_{1}$, Llama-3.1-8B-Instruct[64] as our LLM $L_{2}$. All experiments are run on a single 40G A100 GPU. Additional detailed experiment setup is provided in Appendix A. All of our code is available on Github [^1] https://github.com/AlfredClemedtson/graphraft.
@@
We benchmark our method on the stark-prime and stark-mag datasets[65]. stark-prime is a set of Q&As over a large biomedical knowledge graph PrimeKG[66]. The questions mimic roles such as doctors, medical scientists and patients. It contains 10 node types and 18 edge types with rich textual properties on the nodes. With 129k nodes and 8 million edges and at the same time a high density (average node degree 125), it serves as a challenging and highly suitable dataset to benchmark retrieval and reasoning on large real-world KGs. stark-mag is a set of Q&As over ogbn-mag[67] that models a large academic citation network with relations between papers, authors, subject areas and institutions.
There is a large collection of Q&A datasets on graphs, we benchmark on datasets most suitable to the GraphRAG domain[68] and elaborate on why we don’t benchmark on the others. WebQ[69], WebQSP[4], CWQ[6] ad GrailQA[70] are popular KBQA benchmarks containing SPARQL-answerable few-hop questions over Freebase[5], a database of general knowledge with URIs on nodes. Our problem setting has no requirement on the form of the questions and targets private KGs instead of such largely wikipedia-based KG which LLMs are explicitly trained on. Freebase has been deprecated since 2015. HotpotQA[71] and BeerQA[72] are few-hop questions over Wikidata[73]. STaRK-amazon[65] models properties of products (such as color) as nodes (with has-color relation) and the product co-purchasing graph itself is homogeneous.
@@ 5.1 Main results
We show our result in Table 1. The baselines range from pure vector-based retrievers to GraphRAG solutions to agentic methods. We use the four metrics originally proposed in [65]. Hit@1 measure the ability of exact answering a right answer while the other three metrics provides more holistic view on the answer quality.
As is shown in Table 1, GraphRAFT gives best results on all metrics on both STaRK-prime and STaRK-mag. Even without using $L_{2}$ to reason over the local subgraph, our $L_{1}$ when used to retrieve nodes using generated Cypher queries (up to 20 nodes, for measuring recall), already gives better metrics than all SOTA methods on STaRK-prime and is close to the best baselines for STaRK-mag.
@@
BM25[74] 12.75 27.92 31.25 19.84 25.85 45.25 45.69 34.91
voyage-l2-instruct[75] 10.85 30.23 37.83 19.99 30.06 50.58 50.49 39.66
GritLM-7b[76] 15.57 33.42 39.09 24.11 37.90 56.74 46.40 47.25
multi-ada-002[61] 15.10 33.56 38.05 23.49 25.92 50.43 50.80 36.94
ColBERTv2[77] 11.75 23.85 25.04 17.39 31.18 46.42 43.94 38.39
@@
HybGRAG[78] 28.56 41.38 43.58 34.49 65.40 75.31 65.70 69.80
AvaTaR[79] 18.44 36.73 39.31 26.73 44.36 59.66 50.63 51.15
MoR[80] 36.41 60.01 63.48 46.92 58.19 78.34 75.01 67.14
KAR[81] 30.35 49.30 50.81 39.22 50.47 65.37 60.28 57.51
MFAR[82] 40.9 62.8 68.3 51.2 49.00 69.60 71.79 58.20
GraphRAFT w/o $L_{2}$ 52.12 71.55 75.52 60.72 62.63 71.86 72.97 66.28
GraphRAFT 63.71 75.39 76.39 68.99 71.05 85.34 76.96 76.92
@@ 5.2 Impact of constrained decoding and scaling with training data
We measure the benefit of applying grounded constrained decoding to using our model without it. We also examine how well the method scales with the availability of training data with and without constrained decoding. We measure the metrics directly on the output Cypher queries, executing the ones with highest probabilities first, if executable, and falling back to lower-probability queries, until there are up to 20 nodes. We do not apply LLM $L_{2}$ to reason over the local subgraph to improve Hits@1 to accurately evaluate LLM $L_{1}$ and constrained decoding.
@@
Method (Training data used) Hit@1 Hit@5 Recall@20 MRR
Finetuned, 100% 44.20(43.75) 69.20(63.39) 75.87(69.83) 55.28(52.55)
Finetuned, 10% 41.07(35.71) 66.07(54.91) 76.11(60.53) 51.97(44.00)
LLM, 0% 14.73(0.0) 23.21(0.0) 27.53(0.0) 18.65(0.0)
@@ 5.3 The use of query engine
An alternative valid but suboptimal query plan is provided in Table 4. It starts with scanning all x3:Disease nodes and traverses towards x2 and then x1. It happens to be suboptimal due to the exploding neighbourhood of Disease nodes along the INDICATION relationship type. The query engine with it’s cardinality estimator therefore rules out executing this sequence of operators.
@@
Table 3: The optimal execution plan for an example retrieval query: MATCH (x1:GeneOrProtein {name: "CYP3A4"})-[r: ENZYME]-(x2: Drug)-[r2: %INDICATION]-(x3: Disease {name: "strongyloidiasis"}) RETURN x2.name. Operator are executed from the bottom up. Details represent the exact execution parameters. Estimated Rows represent expected rows produced. Rows represent actual rows produced. DB Hits measure the amount of work by the storage engine. Total database access is 103989, total allocated memory is 328 bytes.
@@
Table 4: A valid but suboptimal query plan for an example retrieval query: MATCH (x1:GeneOrProtein {name: "CYP3A4"})-[r: ENZYME]-(x2: Drug)-[r2: %INDICATION]-(x3: Disease {name: "strongyloidiasis"}) RETURN x2.name. Compared with the optimal plan in Table 3, this plan has both more operators and more costly ones.
@@
5.4 Impact of named entity recognition and ground-truth Cyphers.
Given $\{Q_{i},A_{i}\}$ as input data, we prepare a set of $\{Q_{i},C_{i}\}$ as a set of training data to finetune the LLM to generate optimal Cypher queries. We first few-shot prompt an LLM to identify entities and then reconcile the identified names against the DB. Next, by using the graph schema, we obtain the set of all possible Cypher queries with type predicates k-hop around or connecting the entity nodes. The quality of the identified entities is therefore implicitly measured by the best Cypher that it allows.
We first verify that using an LLM for entity resolution is needed better than simpler methods such as k-nearest-neighbour (kNN) using text embeddings. We measure that by looking at the quality of the best Cypher created from entities identified by an LLM, 2NN and 5NN, as shown in Figure 5.

### v2

@@ 4.3 Finetuning LLM as local subgraph reasoner
Given any question for KG $\mathcal{G}$, our trained LLM $L_{1}$ produces guaranteed executable and grounded Cypher queries. If we perform beam search with width $m$, we are able to obtain $m$ valid queries with highest total probabilities. For the simplest questions and sparse $\mathcal{G}$, the subgraph produced by the best query may be exactly the answer nodes. However, for more difficult multi-hop questions on dense $\mathcal{G}$, the subgraph still contains other relevant nodes (and edges) that are not the answers. In order to obtain only the answer nodes by reasoning over the small subgraph, which often requires reasoning on the textual properties beyond graph patterns, we train an LLM $L_{2}$ to perform the task. We construct the subgraph by executing $m$ queries until some threshold on the size of the graph or tokenization required to encode the graph. The output of $L_{2}$ can be viewed as selecting or reranking the nodes in the input textualised graph. An example prompt we use is provided in Figure 4.
@@ 5 Experiments
We use Neo4j as the graph database. The default database configuration is used. For the main result, we using OpenAI text-embedding-ada-002[openai] as the text embedder $LM$ in Equation 1, OpenAI gpt-4o-mini as LLM $L_{0}$ for few-shot entity resolution., gemma2-9b-text2cypher[ozsoy2024text2cypher, team2024gemma] as our LLM $L_{1}$, Llama-3.1-8B-Instruct[grattafiori2024llama] as our LLM $L_{2}$. All experiments are run on a single 40GB A100 GPU. Additional detailed experiment setup is provided in Appendix B. All of our code is available on Github [^1] https://github.com/AlfredClemedtson/graphraft.
@@
We benchmark our method on the STaRK-prime and STaRK-mag datasets[wu2025stark] (license: CC-BY-4.0). stark-prime is a set of Q&As over a large biomedical knowledge graph PrimeKG[chandak2023building]. The questions mimic roles such as doctors, medical scientists and patients. It contains 10 node types and 18 edge types with rich textual properties on the nodes. With 129k nodes and 8 million edges and at the same time a high density (average node degree 125), it serves as a challenging and highly suitable dataset to benchmark retrieval and reasoning on large real-world KGs. stark-mag is a set of Q&As over ogbn-mag[hu2020open] that models a large academic citation network with relations between papers, authors, subject areas and institutions.
There is a large collection of Q&A datasets on graphs, we benchmark on datasets most suitable to the GraphRAG domain[bechler2025position] and elaborate on why we don’t benchmark on the others. WebQ[Berant2013SemanticPO], WebQSP[Yih2016TheVO], CWQ[talmor-berant-2018-web] ad GrailQA[gu2021beyond] are popular KBQA benchmarks containing SPARQL-answerable few-hop questions over Freebase[freebase], a database of general knowledge with URIs on nodes. Our problem setting has no requirement on the form of the questions and targets private KGs instead of such largely wikipedia-based KG which LLMs are explicitly trained on. Freebase has been deprecated since 2015. HotpotQA[yang2018hotpotqa] and BeerQA[qi2021answering] are few-hop questions over Wikidata[wikidata]. STaRK-amazon[wu2025stark] models properties of products (such as color) as nodes (with has-color relation) and the product co-purchasing graph itself is homogeneous.
@@ 5.1 Main results
We show our result in Table 1. The baselines range from pure vector-based retrievers to GraphRAG solutions to agentic methods. We use the four metrics originally proposed in [wu2025stark]. Hit@1 measure the ability of exact answering a right answer while the other three metrics provides more holistic view on the answer quality.
As is shown in Table 1, GraphRAFT gives best results on all metrics on both STaRK-prime and STaRK-mag. Even without using $L_{2}$ to reason over the local subgraph, our $L_{1}$ when used to retrieve nodes using generated Cypher queries (up to 20 nodes, for measuring recall), already gives better metrics than all SOTA methods on STaRK-prime.
@@
BM25[Robertson2009ThePR] 12.75 27.92 31.25 19.84 25.85 45.25 45.69 34.91
voyage-l2-instruct[voyageai] 10.85 30.23 37.83 19.99 30.06 50.58 50.49 39.66
GritLM-7b[muennighoff2024generative] 15.57 33.42 39.09 24.11 37.90 56.74 46.40 47.25
multi-ada-002[openai] 15.10 33.56 38.05 23.49 25.92 50.43 50.80 36.94
ColBERTv2[santhanam-etal-2022-colbertv2] 11.75 23.85 25.04 17.39 31.18 46.42 43.94 38.39
@@
HybGRAG[lee2024hybgraghybridretrievalaugmentedgeneration] 28.56 41.38 43.58 34.49 65.40 75.31 65.70 69.80
AvaTaR[wu2025avatar] 18.44 36.73 39.31 26.73 44.36 59.66 50.63 51.15
MoR[lei2025mixturestructuralandtextualretrievaltextrich] 36.41 60.01 63.48 46.92 58.19 78.34 75.01 67.14
KAR[xia2025knowledgeawarequeryexpansionlarge] 30.35 49.30 50.81 39.22 50.47 65.37 60.28 57.51
MFAR[li2024multifieldadaptiveretrieval] 40.9 62.8 68.3 51.2 49.00 69.60 71.79 58.20
GraphRAFT w/o $L_{2}$ 52.12 71.55 75.52 60.72 62.63 86.68 88.88 73.03
GraphRAFT 63.71 75.39 76.39 68.99 69.64 84.32 89.12 76.24
@@ 5.2 Impact of constrained decoding and scaling with training data
We measure the benefit of applying grounded constrained decoding to using our model without it. We also examine how well the method scales with the availability of training data with and without constrained decoding. We measure the metrics directly on the output Cypher queries, executing the ones with highest probabilities first, until there are 20 nodes. To accurately evaluate LLM $L_{1}$ and constrained decoding we do not apply LLM $L_{2}$.
@@
Method (% Training data used) Hit@1 Hit@5 Recall@20 MRR
Finetuned, 100% 44.20(43.75) 69.20(63.39) 75.87(69.83) 0.5528(0.5255)
Finetuned, 10% 41.07(35.71) 66.07(54.91) 76.11(60.53) 0.5197(0.4400)
LLM, 0% 14.73(0.0) 23.21(0.0) 27.53(0.0) 0.1865(0.0)
@@ 5.3 The use of query engine
An alternative valid but suboptimal query plan is provided in Appendix C Table 4. It starts with scanning all x3:Disease nodes and traverses towards x2 and then x1. It happens to be suboptimal due to the exploding neighbourhood of Disease nodes along the INDICATION relationship type. The query engine with it’s cardinality estimator therefore rules out executing this sequence of operators.
@@
Table 3: The optimal execution plan for an example retrieval query: MATCH (x1:GeneOrProtein {name: "CYP3A4"})-[r: ENZYME]-(x2: Drug)-[r2: INDICATION]-(x3: Disease {name: "strongyloidiasis"}) RETURN x2.name. Operator are executed from the bottom up. Details represent the exact execution parameters. Estimated Rows represent expected rows produced. Rows represent actual rows produced. DB Hits measure the amount of work by the storage engine. Total database access is 103989, total allocated memory is 328 bytes.
@@
5.4 Ablation study on choices of LLMs
We first verify that using an LLM $L_{0}$ for entity resolution is needed better than simpler methods such as k-nearest-neighbour (kNN) using text embeddings. We measure that by looking at the quality of the best Cypher created from entities identified by an LLM, 2NN and 5NN. Plots showing the quality of entity resolution by various criteria are shown in Appendix D Figure 5.
We also study the different choices of base LLMs for $L_{1}$. In addition to using a general gemma-based text2cypher model, we also finetune a google/gemma2-9b-it and a meta-llama/Llama-3.1-8B-Instruct model. Metrics on STaRK-prime suggest using a already-fintuned general text2cypher model offers some advantage, and gemma2 performs slightly better tham Llama-3.1. Finetuning any of these base models for $L_{1}$ already achieves beyond SOTA results even without $L_{2}$. The table is shown in Appendix D Table 5.
6 Conclusion
In this work, we introduce GraphRAFT, a simple and modular method that leverages graph DBs by retrieving from it using learnt provably correct and optimal Cypher queries. Our experiments show that GraphRAFT consistently achieves beyond SOTA results using smaller LLMs that fit into a single GPU. Our framework can be applied off-the-shelf to any KG in any domain stored in graph DBs. The finetuning process is sample-efficient and scales with more training data.
Appendix A Constrained decoding
We formally state the informal Lemma 1 as three separate formal statements and prove them.
Recall notations in Section 4.2 where $\mathcal{Q}=\{\mathcal{Q}^{1},\cdots,\mathcal{Q}^{M}\}$ is the set of $M$ valid queries for a question, each has tokenization $\mathcal{Q}^{k}=(n_{0},\cdots,n_{Q_{k}})\in\mathcal{V}^{Q_{k}}$ of variable length. We use $l$ and $p$ to denote regular logits and probabilities and $\tilde{l}$ and $\tilde{p}$ to denote those produced by constrained decoding.
Lemma 1.1.
If $Q\notin\mathcal{Q}$ and $Q$ has tokenization $(t_{0},\dots,t_{q})$, then $\tilde{p}(Q)=\tilde{p}(t_{q}|t_{q-1},\dots,t_{0})\times\dots\times\tilde{p}(t_{0})=0$.
Proof.
If $t_{0}$ is not a valid first token, $\mathcal{Q}^{k}_{0}\neq t_{0}$ for all $k\in[1,M]$. Let the logits for the first token produced by LLM $L_{1}$ be $(l_{0},\dots,l_{\mathcal{V}})$. Let the logits after applying constrained decoding be $(\tilde{l}_{0},\dots,\tilde{l}_{\mathcal{V}})$. Then $\tilde{l}_{n_{0}}=$-$\infty$ and the probability of generating token $t_{0}$ is $\tilde{p}(n_{0})=0$.
Now let $(t_{0},\dots,t_{i})$ be the longest subsequence of valid tokens for $Q$. This means there are valid queries $Q^{k_{0}},\dots,Q^{k_{n}}$ such that $Q^{k_{j}}_{0,i}=(t_{0},\dots,t_{i})$ for $j\in[0,n]$. Since this is the longest subsequence, $Q^{k_{j}}_{i+1}\neq t_{i+1}$. For any $j$, let the logits of $i+1^{\text{th}}$ token be $(l_{0},\dots,l_{\mathcal{V}})$ and $(\tilde{l}_{0},\dots,\tilde{l}_{\mathcal{V}})$ after constrained decoding. $\tilde{l}_{t_{i+1}}=$-$\infty$. Hence $\tilde{p}(t_{i+1}|t_{i},\dots,t_{0})=0$. Therefore, $\tilde{p}(Q)=\tilde{p}(t_{q}|t_{q-1},\dots,t_{0})\times\dots\times\tilde{p}(n_{0})=0$ in all cases. ∎
Lemma 1.2.
Suppose that LLM $L_{1}$ has generated a sequence of valid tokens $(t_{0},\dots,t_{k})$, logits for the next token $k+1$ are $(l_{0},\dots,l_{\mathcal{V}})$. If $t_{k+1}$ and $t^{\prime}_{k+1}$ are two valid next tokens and $l_{t_{k+1}}>l_{t^{\prime}_{k+1}}$, then $\tilde{p}(t_{k+1}|t_{k},\dots,t_{0})>\tilde{p}(t^{\prime}_{k+1}|t_{k},\dots,t_{0})$.
Proof.
Since the two next tokens are valid, $l_{t_{k+1}}=\tilde{l}_{t_{k+1}}$ and $l_{t^{\prime}_{k+1}}=\tilde{l}_{t^{\prime}_{k+1}}$. $\tilde{p}(t_{k+1}|t_{k},\dots,t_{0})=\text{softmax}(\tilde{l}_{0},\dots,\tilde{l}_{\mathcal{V}})_{t_{k+1}}=e^{\tilde{l}_{t_{k+1}}}/\sum_{i=0}^{\mathcal{V}}e^{\tilde{l}_{i}}>e^{\tilde{l}_{t^{\prime}_{k+1}}}/\sum_{i=0}^{\mathcal{V}}e^{\tilde{l}_{i}}=\tilde{p}(t^{\prime}_{k+1}|t_{k},\dots,t_{0})$ simply by the monotonicity of the exponential function. ∎
Lemma 1.3.
When beam width = M, LLM $L_{1}$ with constrained decoding generates queries $\{Q^{1},\dots,Q^{M}\}=\{\mathcal{Q}^{1},\dots,\mathcal{Q^{M}}\}$.
Proof.
By the contrapositive of Lemma 1.1, $\tilde{p}(Q)\neq 0$ implies $Q\in\mathcal{Q}$. Hence, any query generated under constrained decoding is a valid query. $\{Q^{1},\dots,Q^{M}\}\subseteq\{\mathcal{Q}^{1},\dots,\mathcal{Q^{M}}\}$. Equality follows from cardinality.
Appendix B Experimental details
All experiments are done on a single machine with one 40GB A100 GPU, 85G RAM 12vCPUs. We use Neo4j Community Edition database with default database configurations. Total time for reproducing results for one datasets (e.g STaRK-prime), from loading raw data into the database, finetuning and evaluting metrics, is around 18 hours.
B.1 Few-shot prompt for entity resolution
B.1.1 STaRK-prime
Answer : "fucose metabolism"
Answer : "A2M gene/protein|PDGFR-beta receptor"
Answer : "melanoma|TNFSF8"
LLM insturction: "You are a knowledgeable assistant which identifies medical entities in the given sentences. Separate entities using '|'."
B.1.2 STaRK-mag
Question : "What research on water absorption in different frequency ranges have been referenced or deemed significant in the paper entitled 'High-resolution terahertz atmospheric water vapor continuum measurements'"
Answer: "FieldOfStudy:water absorption in different frequency ranges
Paper:High-resolution terahertz atmospheric water vapor continuum measurements"
[…]
B.2 K-hop query path templates
MATCH (src {name: srcName})-[r]-(tgt)
and for length-two paths connecting entities, our query is:
UNWIND \$src_names AS srcName1
UNWIND \$src_names AS srcName2
MATCH (src1 {name: srcName1})-[r1]-(tgt)-[r2]-(src2 {name: srcName2})
WHERE src1 <> src2
B.3 Finetuning experimental setups
B.3.1 Training LLM $L_{1}$
For finetuning LLM $L_{1}$ we use huggingface transformer library with LoRA. For base model neo4j/text2cypher-gemma-2-9b-it-finetuned-2024v1 (license: apache-2.0) and google/gemma-2-9b-it (license: gemma), we have
[…]
end_of_generation_token = "<eos>"
bnb_config = BitsAndBytesConfig(load_in_4bit=True, bnb_4bit_use_double_quant=True, bnb_4bit_quant_type="nf4",
bnb_4bit_compute_dtype=torch.bfloat16, )
tokenizer.padding_side = "right"
model = AutoModelForCausalLM.from_pretrained(
model_dir,
quantization_config=bnb_config,
torch_dtype=torch.bfloat16,
attn_implementation="eager",
low_cpu_mem_usage=True,
lora_config = LoraConfig(r=64, lora_alpha=64, target_modules=None, lora_dropout=0.05, bias="none", task_type="CAUSAL_LM", )
For base model meta-llama/Llama-3.1-8B-Instruct (license: llama3.1) used in ablation, the differences are in the tokenization and padding:
[…]
end_of_generation_token = "<|eot_id|>"
tokenizer.pad_token = '<|finetune_right_pad_id|>'
tokenizer.padding_side = "right"
The supervised finetuning training loop is configured as:
sft_config = SFTConfig(auto_find_batch_size=True,
gradient_accumulation_steps=1, #8
dataset_num_proc=8,
num_train_epochs=1,
learning_rate=2e-5,
optim="paged_adamw_8bit",
max_seq_length=max_seq_len,
eval_strategy="epoch",
save_strategy="epoch",
logging_steps=10,
output_dir=model_save_dir,
load_best_model_at_end=True,
B.3.2 Training LLM $L_{2}$
For finetuning LLM $L_{2}$ we use unsloth[unsloth]. For the base model meta-llama/Llama-3.1-8B-Instruct (license: llama3.1), our configurations are:
[…]
EOS = "<|eot_id|>"
RIGHT_PAD_TOKEN = '<|finetune_right_pad_id|>'
ANSWER_SEPARATOR = '|'
PRIME_MAX_SEQUENCE_LENGTH = 15_000
MAG_MAX_SEQUENCE_LENGTH = 15_000
MAX_NEW_TOKENS = 100
[…]
model, tokenizer = FastLanguageModel.from_pretrained(
model_name=model_dir,
max_seq_length=max_sequence_length, # None
dtype=torch.bfloat16,
load_in_4bit=True,
tokenizer.padding = True
tokenizer.pad_token = RIGHT_PAD_TOKEN
tokenizer.padding_side = 'right'
model = FastLanguageModel.get_peft_model(
model=model,
r=64,
target_modules=["q_proj", "k_proj", "v_proj", "o_proj", "gate_proj", "up_proj", "down_proj", ],
lora_alpha=64,
bias="none",
sft_config = SFTConfig(
per_device_train_batch_size=1,
per_device_eval_batch_size=1,
# auto_find_batch_size=True,
dataset_num_proc=8,
bf16=True,
num_train_epochs=1,
gradient_accumulation_steps=4,
warmup_steps=5,
learning_rate=2e-5,
optim="adamw_8bit",
weight_decay=0.01,
lr_scheduler_type="linear",
logging_steps=1,
eval_strategy="epoch",
save_strategy="epoch",
output_dir=model_save_dir,
load_best_model_at_end=True,
Appendix C Analysis of alternative query plans
An alternative valid but suboptimal query plan for the same Cypher query in Table 3 is shown in Table 4. The query optimiser rules out executing this less optimal plan. Adhoc retrieval methods in GraphRAG that do not use the underlying databases either require essentially a rewrite of the query engine in-memory, or risks having suboptimal retrievals.
@@ Appendix C Analysis of alternative query plans
Table 4: A valid but suboptimal query plan for an example retrieval query: MATCH (x1:GeneOrProtein {name: "CYP3A4"})-[r: ENZYME]-(x2: Drug)-[r2: INDICATION]-(x3: Disease {name: "strongyloidiasis"}) RETURN x2.name. Compared with the optimal plan in Table 3, this plan has both more operators and more costly ones.
@@
Appendix D Ablation of choices of LLMs
Given $\{Q_{i},A_{i}\}$ as input data, we prepare a set of $\{Q_{i},C_{i}\}$ as a set of training data to finetune the LLM to generate optimal Cypher queries. We first few-shot prompt an LLM to identify entities and then reconcile the identified names against the DB. Next, by using the graph schema, we obtain the set of all possible Cypher queries with type predicates k-hop around or connecting the entity nodes. The quality of the identified entities is therefore implicitly measured by the best Cypher that it allows. Figure 5 shows that using an LLM for entity resolution does produce better entities that lead to more high-recall ground-truth Cypher queries with more accurate subgraphs returned.